# On regions of mixed unitarity for semigroups of unital quantum channels

### v1

@@ 4. Quantum dynamical semigroups of unital quantum channels
$(iii)$ For every $t\geq 0$, let $\alpha_{t}:=e^{t\mathcal{L}}\circ\mathbf{E}_{\mathcal{P}}$ and $\beta_{t}:=e^{t\mathcal{L}}-\alpha_{t}$. Given $Z\in\mathbb{M}_{d}$, decompose it as $Z=X+Y$ with $X\in\mathcal{P}$ and $Y\in\mathcal{N}$. Then for all $t\geq 0$,
$\beta_{t}(Z)=e^{t\mathcal{L}}(Z)-\alpha_{t}(Z)=e^{t\mathcal{L}}(X)+e^{t\mathcal{L}}(Y)-e^{t\mathcal{L}}\circ\mathbf{E}_{\mathcal{P}}(Z)=e^{t\mathcal{L}}(Y).$
@@
Since $e^{t\mathcal{L}}(Y)\to 0$ as $t\to\infty$, we have $\beta_{t}(\mathbb{M}_{d})\subseteq\mathcal{N}$ for all $t\geq 0$ and $\left\lVert e^{t\mathcal{L}}-\alpha_{t}\right\rVert=\left\lVert\beta_{t}\right\rVert\to 0$ as $t\to\infty$. ∎
We note that part of (i) of the previous theorem does not require $\Phi_{t}$ to be trace-preserving.
@@
Let $\{\Phi_{t}\}_{t\geq 0}$ be a quantum dynamical semigroup of unital channels on $\mathbb{M}_{d}$. We use the notation and results of Theorem 4.6. Now, from Theorem 3.6, there exists unitaries $\{U_{t}\}_{t\geq 0}\subseteq\mathbb{U}(d)$ such that
@@
Take $\tau_{t}:=\mathrm{Ad}_{U_{t}^{*}}\circ\Phi_{t}$. Then for every $X\in\mathcal{P}$,
$\tau_{t}(X)$ $=\mathrm{Ad}_{U_{t}^{*}}\circ\Phi_{t}(X)=\mathrm{Ad}_{U_{t}^{*}}\circ\mathrm{Ad}_{U_{t}}(X)=X.$
Note that, $\mathbf{E}_{\mathcal{P}}=\mathrm{Ad}_{U_{t}^{*}}\circ\alpha_{t}$ for all $t\geq 0$, therefore
@@
Let $r_{\mathcal{P}}>0$ be as in the Theorem 3.4. Since $\beta_{t}\to 0$ as $t\to\infty$, there exists a $t_{0}\geq 0$ such that $\left\lVert\tau_{t}-\mathbf{E}_{\mathcal{P}}\right\rVert\leq r_{\mathcal{P}}$, for all $t\geq t_{0}$. From Theorem 3.4 (ii), it then follows that $\tau_{t}$ is mixed unitary for all $t\geq t_{0}$. Consequently, $\Phi_{t}=\mathrm{Ad}_{U_{t}}\circ\tau_{t}$ is mixed unitary for all $t\geq t_{0}$. ∎
@@
[HRS20] For any unital trace-preserving linear map $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$, the mixed unitary time or mixed unitary index of $\Phi$ is defined as the least $n\in\mathbb{N}$ such that $\Phi^{k}$ is mixed unitary for all $k\geq n$.
From Theorem 3.9, it follows that the mixed unitary index of a unital quantum channel is finite. One can ask whether the mixed unitary index is independent of $\Phi$. i.e., Is there an integer $n$ such that for every unital quantum channel $\Phi$ on $\mathbb{M}_{d}$ we have $\Phi^{k}$ is mixed unitary, $\forall k\geq n$. The answer is negative. That is, there is no universal power $n$ such that $\Phi^{n}$ mixed unitary for all unital quantum channel $\Phi$ on $\mathbb{M}_{d}$.
@@
Suppose that there exists $n\geq 2$ such that $\Phi^{n}$ is mixed unitary for all unital quantum channels $\Phi$ on $\mathbb{M}_{d}$. Let $\mathcal{L}$ be generator of a quantum dynamical semigroup of unital channels, which is not mixed unitary for some $t_{0}>0$. (Note that such $\mathcal{L}$ exists. For example, take $\mathcal{L}$ as in the Lemma 5.9.) Consider $\Phi:=e^{\frac{t_{0}}{n}\mathcal{L}}$. Then $\Phi$ is a unital quantum channel on $\mathbb{M}_{d}$, but $\Phi^{n}=e^{t_{0}\mathcal{L}}$ is not mixed unitary, which gives a contradiction. ∎
@@ 5. Behavior of dynamical semigroup of quantum channels near origin
In other words, $\mathscr{W}(d)$ is the set of unital, trace-preserving Hermitian maps on $\mathbb{M}_{d}$ and $\mathscr{W}_{1}(d)$ is the collection of unit elements in it. Further, $\mathscr{D}(d)$ is a convenient subset of the dual of the cone generated by mixed unitary channels. Note that the inner product here is defined using Choi matrices as in 2.2. This will be used in the characterization below in Theorem 5.2.
@@
$0\leq\langle\Gamma,\operatorname{\textsf{id}}\rangle=\lim_{n_{k}}\langle\Gamma_{n_{k}},\operatorname{\textsf{id}}\rangle<\lim_{n_{k}}\frac{1}{n_{k}}=0.$
@@
Recall that Kümmerer et al. [KuMa87] characterizes (see Theorem 4.3) the dynamical semigroups that are mixed unitary channels for all $t\geq 0$. Here, we provide another necessary and sufficient condition for a semigroup to be mixed unitary.
@@
Therefore, for all $\Phi\in\text{cone}\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in\mathbb{U}(d)\}$ we have $\langle\Gamma,\Phi\rangle\geq 0$. Consequently we also get $\langle\Gamma,\Phi\rangle\geq 0$, for all $\Phi$ is in the closed cone generated by $\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in\mathbb{U}(d)\}$. In particular, we get $\langle\Gamma,\mathcal{L}\rangle\geq 0$.
@@
    There exist $0<t_{0}\leq t_{1}<\infty$ such that $\Phi_{t}$ is not mixed unitary for $t\in(0,t_{0})$ and is mixed unitary for $t\in[t_{1},\infty)$.
@@
We wish to remark that in the previous theorem we do not know whether $t_{0}<t_{1}$ is feasible or not. Considering the general features of numerous situations of eventual positivity (See for example [Ja17]) we would guess that this could be possible for some examples with the set of points in $t_{0},t_{1}]$ for which $\Phi_{t}$ is a mixed unitary channel forming a discrete set. We are not able to confirm this in any of the following examples.
@@
[MeWo09, Corollory 15] Given non-negative numbers $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{d}$ we have
@@
$=\frac{1}{18}\min\{\operatorname{\textsf{tr}}(A\overline{A}):\mu(A)=(1,1,1)\}+\frac{1}{18}$ $(\because\mu(B)=(1,1,1))$
@@
[KuMa87] Let $A\in\mathbb{M}_{d}$ be such that either $A$ is non-normal, or its spectrum is neither contained in a circle nor a straight line. Take the linear map $\mathcal{L}$ on $\mathbb{M}_{d}$
@@
where $B=\begin{bmatrix}1&0&0\\ 0&0&-1\\ 0&1&0\end{bmatrix}$. Consider the semigroup $e^{t\mathcal{L}}$. Then, by Theorem 4.2 $e^{t\mathcal{L}}$ is a dynamical semigroup of unital quantum channels. Let $t_{0}\approx 1.4034$ be the positive root of the function $e^{t}-e^{-\frac{t}{2}}-3\sinh(\frac{t}{2})-2\sin(\frac{t}{2})$.
@@
Proof of the claim : For $1\leq j\leq 3$, let $\mathcal{L}_{j}$ be the linear maps on $\mathbb{M}_{3}$ given by
@@
Then $\mathcal{L}_{j}$’s commute with each other, and
@@
To compute $e^{t\mathcal{L}_{1}}$, first note that $\mathrm{Ad}_{B^{\operatorname{\textsf{T}}}}\circ\operatorname{\textsf{T}}=\mathrm{Ad}_{B^{*}}\circ\operatorname{\textsf{T}}=\operatorname{\textsf{T}}\circ\mathrm{Ad}_{B^{\operatorname{\textsf{T}}}}$, $B^{2}=(B^{*})^{2}=(B^{\operatorname{\textsf{T}}})^{2}$ and $B=(B^{*})^{3}=(B^{\operatorname{\textsf{T}}})^{3}$,
@@
Furthermore, we have
@@
$=\frac{1}{2}\big(\operatorname{\textsf{tr}}(\mathscr{J}(\operatorname{\textsf{id}})\mathscr{J}(\mathrm{Ad}_{B}\circ\operatorname{\textsf{T}}))+\operatorname{\textsf{tr}}(\mathscr{J}(\operatorname{\textsf{id}})\mathscr{J}(\delta_{3}))=\frac{1}{6}\big(\langle\Omega_{d},(I_{3}\otimes B^{*})F(I_{3}\otimes B)\Omega_{d}\rangle+\frac{1}{3}\langle\Omega_{d},\Omega_{d}\rangle\big)$

### v2

@@ 4. Quantum dynamical semigroups of unital quantum channels
Proof.
Suppose $X\in\mathcal{D}=\bigcap_{t\geq 0}\mathscr{M}_{\Phi_{t}}$ then
$\Phi_{t+s}(X^{*}X)=\Phi_{t}\big(\Phi_{s}(X^{*}X)\big)$ $=\Phi_{t}\big(\Phi_{s}(X^{*})\Phi_{s}(X)\big)$
$\geq\Phi_{t}\big(\Phi_{s}(X^{*})\big)\Phi_{t}\big(\Phi_{s}(X)\big)$
$=\Phi_{t+s}(X^{*})\Phi_{t+s}(X)=\Phi_{t+s}(X^{*}X),\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall s,t\geq 0.$
This shows that $\Phi_{t}\big(\Phi_{s}(X^{*})\Phi_{s}(X)\big)=\Phi_{t}\big(\Phi_{s}(X^{*})\big)\Phi_{t}\big(\Phi_{s}(X)\big)$ and consequently, we get $\Phi_{s}(X)\in\mathcal{D}$, for all $s\geq 0$. Since $\mathcal{D}\subseteq\mathscr{M}_{\Phi_{s}}$, $\Phi_{s}$ is a *-homomorphism on $\mathcal{D}$, for all $s\geq 0$. Now, to prove that it is bijective it is enough to show that $\Phi_{s}$ is injective. For this, let $X\in\mathcal{D}$ such that $\Phi_{s}(X)=0$ then
$\langle X,X\rangle=\operatorname{\textsf{tr}}(X^{*}X)=\operatorname{\textsf{tr}}\big(\Phi_{s}(X^{*}X)\big)=\operatorname{\textsf{tr}}\big(\Phi_{s}(X^{*})\Phi_{s}(X)\big)=0.$
Hence $X=0.$ This completes the proof. ∎
Theorem 4.8.
Let $\{\Phi_{t}=e^{t\mathcal{L}}\}_{t\geq 0}$ be a quantum dynamical semigroup of unital channels on $\mathbb{M}_{d}$ with $\mathcal{L}$ as in (4.1). Then, the following spaces are equal:
  • The peripheral Poisson boundary $\mathcal{P}:=\operatorname{\textsf{span}}\{X\in\mathbb{M}_{d}:\mathcal{L}(X)=iaX,\text{ for some }a\in\mathbb{R}\}.$
  • The decoherence-free algebra $\mathcal{D}:=\bigcap_{t\geq 0}\mathscr{M}_{\Phi_{t}}$.
  • $\mathcal{C}:=\operatorname{\textsf{span}}\{X\in\mathbb{M}_{d}:i[H,X]=iaX,[L_{j},X]=[L_{j}^{*},X]=0,1\leq j\leq n,\text{ for some }a\in\mathbb{R}\}.$
Proof.
By virtue of Theorem 4.6, to show that $\mathcal{D}=\mathcal{P}$, it is enough to prove that $\mathcal{D}^{\perp}=\{Y\in\mathbb{M}_{d}:\lim_{t\to\infty}\Phi_{t}(Y)=0\}$. Suppose $Y\in\mathbb{M}_{d}$ satisfies $\lim_{t\to\infty}\Phi_{t}(Y)=0$. Then for any $X\in\mathcal{D}$, any $t\geq 0$, we have
$\left\lvert\langle X,Y\rangle\right\rvert=\left\lvert\operatorname{\textsf{tr}}(X^{*}Y)\right\rvert=\left\lvert\operatorname{\textsf{tr}}(\Phi_{t}(X^{*}Y))\right\rvert=\left\lvert\operatorname{\textsf{tr}}(\Phi_{t}(X)^{*}\Phi_{t}(Y))\right\rvert\leq d\left\lVert\Phi_{t}(X)\right\rVert\left\lVert\Phi_{t}(Y)\right\rVert\leq d\left\lVert X\right\rVert\left\lVert\Phi_{t}(Y))\right\rVert$
This shows that $\langle X,Y\rangle=0$ and hence $Y\in\mathcal{D}^{\perp}$. Now, to prove the reverse inclusion, let $Y\in\mathcal{D}^{\perp}$. As the set $\{\Phi_{t}(Y)\}_{t\geq 0}$ is bounded, there exists $\{t_{n}\}\subset[0,\infty)$, converging to infinity such that $\lim_{n\to\infty}\Phi_{t_{n}}(Y)=Y_{0},$ for some $Y_{0}\in\mathbb{M}_{d}$. Note that as $\Phi_{t}$ is contractive in Hilbert-Schmidt norm, $\{\left\lVert\Phi_{t}(Y)\right\rVert_{2}^{2}\}_{t\geq 0}$ is non-increasing. Consequently $\lim_{n\to\infty}\|\Phi_{s+t_{n}}(Y)\|_{2}^{2}$ $\|Y_{0}\|_{2}^{2}$ for every $s$. Now, by Kadison-Schwarz inequality, $\Phi_{s}(Y_{0}^{*}Y_{0})-\Phi_{s}(Y_{0}^{*})\Phi_{s}(Y_{0})\geq 0$, and
$0=\lim_{n\to\infty}(\left\lVert\Phi_{t_{n}}(Y)\right\rVert_{2}^{2}-\left\lVert\Phi_{s+t_{n}}(Y)\right\rVert_{2}^{2})$ $=\lim_{n\to\infty}\big(\operatorname{\textsf{tr}}(\Phi_{t_{n}}(Y^{*})\Phi_{t_{n}}(Y))-\operatorname{\textsf{tr}}(\Phi_{s+t_{n}}(Y^{*})\Phi_{s+t_{n}}(Y))\big)$
$=\lim_{n\to\infty}\Big(\operatorname{\textsf{tr}}\big(\Phi_{s}(\Phi_{t_{n}}(Y^{*})\Phi_{t_{n}}(Y))\big)-\operatorname{\textsf{tr}}(\Phi_{s+t_{n}}(Y^{*})\Phi_{s+t_{n}}(Y))\Big)$
$=\operatorname{\textsf{tr}}\big(\Phi_{s}(Y_{0}^{*}Y_{0})-\Phi_{s}(Y_{0}^{*})\Phi_{s}(Y_{0})\big).$
So we get $\Phi_{s}(Y_{0}^{*}Y_{0})=\Phi_{s}(Y_{0}^{*})\Phi_{s}(Y_{0})$, for all $s\geq 0$. Similarly, one can show that $\Phi_{s}(Y_{0}Y_{0}^{*})=\Phi_{s}(Y_{0})\Phi_{s}(Y_{0}^{*})$, for all $s\geq 0$. Therefore, $Y_{0}\in\mathcal{D}$. Now, we will also show that $Y_{0}\in\mathcal{D}^{\perp}$. To see this first observe that from Lemma 4.7, $\Phi_{t}$ restricted to $\mathcal{D}$ is bijective. Given $t\geq 0$ and $Z\in\mathcal{D}$, let $X\in\mathcal{D}$ be the unique element such that $\Phi_{t}(X)=Z$, then
$\langle Z,\Phi_{t}(Y)\rangle=\langle\Phi_{t}(X),\Phi_{t}(Y)\rangle=\operatorname{\textsf{tr}}(\Phi_{t}(X^{*})\Phi_{t}(Y))=\operatorname{\textsf{tr}}(\Phi_{t}(X^{*}Y))=\langle X,Y\rangle=0.$
This shows that $\Phi_{t}(Y)\in\mathcal{D}^{\perp}$ for every $t\geq 0$ and hence $Y_{0}=\lim_{n\to\infty}\Phi_{t_{n}}(Y)\in\mathcal{D}^{\perp}$. Therefore $Y_{0}=0$. Now, since $\{\left\lVert\Phi_{t}\right\rVert_{2}\}_{t\geq 0}$ is non-increasing and $\lim_{n\to\infty}\Phi_{t_{n}}(Y)=0$, we conclude that $\lim_{t\to\infty}\Phi_{t}(Y)=0$. This completes the proof of $\mathcal{D}=\mathcal{P}$.
Observe that for any $Z\in\mathbb{M}_{d}$, by direct computation:
$\mathcal{L}(Z^{*}Z)=Z^{*}\mathcal{L}(Z)+\mathcal{L}(Z^{*})Z+\sum_{j=1}^{n}[L_{j},Z]^{*}[L_{j},Z].$ (4.2)
Suppose $X\in\mathcal{P}$ is such that $\mathcal{L}(X)=iaX$ with $a\in\mathbb{R}$. Then $\Phi_{t}(X)=e^{ita}X$. Since we have assumed that $\Phi_{t}$ is trace preserving, by Kadison-Schwarz inequality, we conclude that $\Phi_{t}(X^{*}X)=X^{*}X$ and $\Phi_{t}(XX^{*})=XX^{*}$, for all $t\geq 0$. By differentiation at $t=0$, we get $\mathcal{L}(X^{*}X)=0=\mathcal{L}(XX^{*})$. Thus, by letting $Z=X$ in equation (4.2), we get $\sum_{j=1}^{n}[L_{j},X]^{*}[L_{j},X]=0$, and hence $L_{j},X]=0$ for every $j$. Similarly by taking taking $Z=X^{*}$, we get $L_{j},X^{*}]=0$ for every $j$. Hence, $iaX=\mathcal{L}(X)=i[H,X]$, proving $\mathcal{P}\subseteq\mathcal{C}.$ The reverse inclusion is clear from the formula (4.1) of $\mathcal{L}$. ∎
The following theorem is the continuous analogue of Theorem 3.8. The conditional automorphism under consideration may not be unique. It depends upon the self-adjoint operator $H$ chosen in (4.1).
Theorem 4.9.
Let $\{\Phi_{t}=e^{t\mathcal{L}}\}$ be a dynamical semigroup of unital channels on $\mathbb{M}_{d}$ with the generator $\mathcal{L}$ given by (4.1). That is,
$\mathcal{L}(X)=i[H,X]-\frac{1}{2}\sum_{j=1}^{n}(L_{j}^{*}L_{j}X+XL_{j}^{*}L_{j}-2L_{j}^{*}XL_{j}),\penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{d},$
for some $H,L_{j},1\leq j\leq n$ in $\mathbb{M}_{d},\penalty 10000\ n\geq 0,$ satisfying $H=H^{*},\penalty 10000\ \sum_{j=1}^{n}L_{j}L_{j}^{*}=\sum_{j=1}^{n}L_{j}^{*}L_{j}.$ Then
@@
$\lim_{t\to\infty}\left\lVert\Phi_{t}-\mathrm{Ad}_{U_{t}}\circ\mathbf{E}_{\mathcal{P}}\right\rVert=0,$
where $U_{t}=e^{-itH},\penalty 10000\ \penalty 10000\ \forall t\geq 0$ and $\mathbf{E}_{\mathcal{P}}$ is the conditional expectation map onto the peripheral Poisson boundary $\mathcal{P}$.
Proof.
Suppose $X$ is such that $i[H,X]=iaX$ for some $a\in\mathbb{R}$ and $L_{j},X]=[L_{j}^{*},X]=0$ for all $j$. Then clearly, $\mathcal{L}(X)=iaX.$ Consequently, by taking the power series expansion of the exponential, we get $e^{ita}X=e^{t\mathcal{L}}(X)=\mathrm{Ad}_{U_{t}}(X)$. Making use of Theorem 4.8, we conclude that $\Phi_{t}(X)=\mathrm{Ad}_{U_{t}}(X)$, for all $X\in\mathcal{P}$. Then for any $Z=X+Y$ with $X\in\mathcal{P}$ and $Y\in\mathcal{P}^{\perp}$, by Theorem 4.6, we have
$\big(\Phi_{t}-\mathrm{Ad}_{U_{t}}\circ\mathbf{E}_{\mathcal{P}}\big)(Z)=\Phi_{t}(X)+\Phi_{t}(Y)-\mathrm{Ad}_{U_{t}}(X)=\Phi_{t}(Y)\to 0,\penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{as }t\to\infty.$
Corollary 4.10.
Let $\{\Phi_{t}=e^{t\mathcal{L}}\}$ be a dynamical semigroup of unital channels on $\mathbb{M}_{d}$ with the generator $\mathcal{L}$ as in (4.1). If $H$ is a scalar matrix, then $\Phi_{t}$ converges to the conditional expectation $\mathbf{E}_{\mathcal{P}}$ as $t$ goes to infinity.
Proof.
Assume that $H$ is a scalar matrix and $U_{t}=e^{-itH}$, then $\mathrm{Ad}_{U_{t}}(X)=X$, for all $X\in\mathbb{M}_{d}$ and thus from Theorem 4.9, $\Phi_{t}$ converges to a conditional expectation onto the peripheral space $\mathcal{P}$. ∎
The converse of this result is not true as the following remark shows.
Remark 4.11.
A dynamical semigroup of unital channels $\{\Phi_{t}\}_{t\geq 0}$ converges to $\mathbf{E}_{\mathcal{P}}$, as $t\to\infty$ does not imply that $H$ is a scalar matrix. For example, take $\mathcal{L}(X)=i[H,X]+\frac{1}{2}(2L^{*}XL-LL^{*}X-XLL^{*})$, where $L$ is a normal matrix such that $L+L^{*}$ is not a scalar matrix and $H=L+L^{*}$. Note that, if $X\in\mathcal{P}$, using the same argument as in Theorem 4.9, we conclude that $L_{j},X]=0=[L_{j}^{*},X]$ and consequently $\mathcal{L}(X)=i[H,X]=i[L_{j}+L_{j}^{*},X]=0$. Now, by taking the power series, we get $\Phi_{t}(X)=X$ for all $t\geq 0$. By writing any $Z=X+Y$ with $X\in\mathcal{P}$ and $Y\in\mathcal{N}$, we have
$\Phi_{t}(Z)=\Phi_{t}(X)+\Phi_{t}(Y)=X+\Phi_{t}(Y)=\mathbf{E}_{\mathcal{P}}(Z)+\Phi_{t}(Y)\to\mathbf{E}_{\mathcal{P}}(Z),\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mbox{as }t\to\infty.$
@@
Let $\{\Phi_{t}\}_{t\geq 0}$ be a quantum dynamical semigroup of unital channels on $\mathbb{M}_{d}$. We use the notation and results of Theorem 4.6 and Theorem 4.9. From the proof of Theorem 4.9, we know that
@@
where $U_{t}=e^{-itH}$. Take $\tau_{t}:=\mathrm{Ad}_{U_{t}^{*}}\circ\Phi_{t}$, then
$\tau_{t}(X)$ $=\mathrm{Ad}_{U_{t}^{*}}\circ\Phi_{t}(X)=\mathrm{Ad}_{U_{t}^{*}}\circ\mathrm{Ad}_{U_{t}}(X)=X,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall X\in\mathcal{P}.$
From Theorem 4.6, we also know that $\Phi_{t}=\alpha_{t}+\beta_{t}$, where $\alpha_{t}=\Phi_{t}\circ\mathbf{E}_{\mathcal{P}}$ and $\lim_{t\to\infty}\beta_{t}=0$. Therefore, $\mathrm{Ad}_{U_{t}^{*}}\circ\alpha_{t}=\mathbf{E}_{\mathcal{P}}$ for all $t\geq 0$, and
@@
By choosing $r_{\mathcal{P}}>0$ as in Theorem 3.4, there exists a $t_{0}\geq 0$ such that $\left\lVert\tau_{t}-\mathbf{E}_{\mathcal{P}}\right\rVert\leq r_{\mathcal{P}}$, for all $t\geq t_{0}$. Now, by Theorem 3.4, it follows that $\tau_{t}$ is mixed unitary for all $t\geq t_{0}$ and consequently, $\Phi_{t}=\mathrm{Ad}_{U_{t}}\circ\tau_{t}$ is mixed unitary for all $t\geq t_{0}$. ∎
We now briefly return to the case of discrete semigroups. The proof of the following Remark is very similar to the proof of Theorem 4.8 and so we omit it.
Remark 4.13.
Let $\Phi$ be a unital channel on $\mathbb{M}_{d}$. Consider the discrete semigroup of unital channels $\{\Phi^{n}:n\geq 1\}$, then its decoherence free algebra is same as the peripheral Poisson boundary. That is, $\bigcap_{n\geq 1}\mathscr{M}_{\Phi^{n}}=\operatorname{\textsf{span}}\{X:\Phi(X)=\lambda X,\penalty 10000\ \left\lvert\lambda\right\rvert=1\}.$
@@
[17] For any unital trace-preserving linear map $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$, the mixed unitary time or mixed unitary index of $\Phi$ is defined as the least $n\in\mathbb{N}$ such that $\Phi^{k}$ is mixed unitary for all $k\geq n$.
From Theorem 3.9, it follows that the mixed unitary index of a unital quantum channel is finite. One can ask whether the mixed unitary index is independent of $\Phi$. i.e., Is there an integer $n$ such that for every unital quantum channel $\Phi$ on $\mathbb{M}_{d}$, $\Phi^{k}$ is mixed unitary, $\forall k\geq n$. The answer is negative as the following result shows.
@@
Suppose that there exists $n\geq 2$ such that $\Phi^{n}$ is mixed unitary for all unital quantum channels $\Phi$ on $\mathbb{M}_{d}$. Let $\mathcal{L}$ be the generator of a quantum dynamical semigroup of unital channels, which is not mixed unitary for some $t_{0}>0$. (Note that such $\mathcal{L}$ exists. For example, take $\mathcal{L}$ as in Lemma 5.9.) For $n\geq 1$, take $\Phi_{n}:=e^{\frac{t_{0}}{n}\mathcal{L}}$. Then $\Phi_{n}$ is a unital quantum channel on $\mathbb{M}_{d}$, but $(\Phi_{n})^{n}=e^{t_{0}\mathcal{L}}$ is not mixed unitary. ∎
@@ 5. Behavior of dynamical semigroup of quantum channels near origin
In other words, $\mathscr{W}(d)$ is the set of unital, trace-preserving Hermitian preserving maps on $\mathbb{M}_{d}$ and $\mathscr{W}_{1}(d)$ is the collection of unit elements in it. Further, $\mathscr{D}(d)$ is a convenient subset of the dual of the cone generated by mixed unitary channels. Note that the inner product here is defined using Choi matrices as in (2.2). This will be used in the characterization below in Theorem 5.2.
@@
$0\leq\langle\Gamma,\operatorname{\textsf{id}}\rangle=\lim_{n_{k}}\langle\Gamma_{n_{k}},\operatorname{\textsf{id}}\rangle\leq\lim_{n_{k}}\frac{1}{n_{k}}=0.$
@@
Recall that Kümmerer et al. [25] characterizes (See Theorem 4.3) the dynamical semigroups that are mixed unitary channels for all $t\geq 0$. Here, we provide another necessary and sufficient condition for a semigroup to be mixed unitary.
@@
Therefore, for all $\Phi\in\text{cone}\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in\mathbb{U}(d)\}$ we have $\langle\Gamma,\Phi\rangle\geq 0$. Consequently, we also get $\langle\Gamma,\Phi\rangle\geq 0$, for all $\Phi$ is in the closed cone generated by $\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in\mathbb{U}(d)\}$. In particular, we get $\langle\Gamma,\mathcal{L}\rangle\geq 0$.
@@
    There exist $0<t_{0}\leq t_{1}<\infty$ such that $\Phi_{t}$ is not mixed unitary for $t\in(0,t_{0})$ and is mixed unitary for $t=t_{0}$ and for $t\in[t_{1},\infty)$.
@@
We wish to remark that in the previous theorem, we do not know whether one can take $t_{1}=t_{0}$ or not. Considering the general features of numerous situations of eventual positivity (See for example [19]) we would guess that there could be examples for which $t_{0}<t_{1}$ and the set of points $t$ in $t_{0},t_{1}]$ for which $\Phi_{t}$ is a mixed unitary channel is a discrete set. We are not able to confirm this in any of the following examples.
@@
[29, Corollory 15] Given non-negative numbers $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{d}$ we have
@@
$=\frac{1}{18}\min\{\operatorname{\textsf{tr}}(A\overline{A}):\mu(A)=(1,1,1)\}+\frac{1}{18}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ (\because\mu(B)=(1,1,1))$
@@
[25] Let $A\in\mathbb{M}_{d}$ be a normal matrix whose spectrum is neither contained in a circle nor on a straight line. Take the linear map $\mathcal{L}$ on $\mathbb{M}_{d}$
@@
where $B=\begin{bmatrix}1&0&0\\ 0&0&-1\\ 0&1&0\end{bmatrix}$. Consider the semigroup $e^{t\mathcal{L}}$. According to Theorem 4.2, this family constitutes a dynamical semigroup of unital quantum channels. Let $t_{0}\approx 1.4034$ be the positive root of the function $e^{t}-e^{-\frac{t}{2}}-3\sinh(\frac{t}{2})-2\sin(\frac{t}{2})$.
@@
Proof of the claim : Write $\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{2}+\mathcal{L}_{3}$, where $\mathcal{L}_{j}$ be the linear maps on $\mathbb{M}_{3}$ given by
@@
Then $\mathcal{L}_{j}$’s commute with each other, and hence $e^{t\mathcal{L}}=e^{t\mathcal{L}_{1}}\circ e^{t\mathcal{L}_{2}}\circ e^{t\mathcal{L}_{3}}$. Consider,
@@
Now to compute $e^{t\mathcal{L}_{1}}$, first note that $\mathrm{Ad}_{B^{\operatorname{\textsf{T}}}}\circ\operatorname{\textsf{T}}=\mathrm{Ad}_{B^{*}}\circ\operatorname{\textsf{T}}=\operatorname{\textsf{T}}\circ\mathrm{Ad}_{B^{\operatorname{\textsf{T}}}}$, $B^{2}=(B^{*})^{2}=(B^{\operatorname{\textsf{T}}})^{2},B=(B^{*})^{3}=(B^{\operatorname{\textsf{T}}})^{3}$ and $BB^{\operatorname{\textsf{T}}}=I_{3}=(B^{\operatorname{\textsf{T}}})^{4}$. This gives us,
@@
and consequently,
$e^{t\mathcal{L}}=(1-e^{-\frac{3}{2}t})\delta_{3}+e^{-t}(a_{t}\operatorname{\textsf{id}}+b_{t}\mathrm{Ad}_{B^{\operatorname{\textsf{T}}}}\circ\operatorname{\textsf{T}}+c_{t}\mathrm{Ad}_{(B^{\operatorname{\textsf{T}}})^{2}}+d_{t}\mathrm{Ad}_{B}\circ\operatorname{\textsf{T}}),$
@@
Let $\Gamma$ be the map given in the Lemma 5.7. That is, $\Gamma=\frac{1}{2}(\mathrm{Ad}_{B^{*}}\circ\operatorname{\textsf{T}}+\delta_{3})$, then we have
@@
$=\frac{1}{2}\big(\operatorname{\textsf{tr}}(\mathscr{J}(\operatorname{\textsf{id}})\mathscr{J}(\mathrm{Ad}_{B}\circ\operatorname{\textsf{T}}))+\operatorname{\textsf{tr}}(\mathscr{J}(\operatorname{\textsf{id}})\mathscr{J}(\delta_{3}))$
$=\frac{1}{6}\big(\langle\Omega_{3},(I_{3}\otimes B^{*})F(I_{3}\otimes B)\Omega_{3}\rangle+\frac{1}{3}\langle\Omega_{3},\Omega_{3}\rangle\big)\qquad(\text{here }F=\sum_{i,j=1}^{3}E_{ij}\otimes E_{ji})$